# RNA-FrameFlow: Flow Matching for de novo 3D RNA Backbone Design

### v3

@@ 1 Introduction
Our Contributions. We develop RNA-FrameFlow, the first generative model for 3D all-atom RNA backbone design, which trains on RNA of length 40-150 from the PDB; see Figure 1. We devise RNA frames to capture the significantly larger RNA nucleotides while still supporting all-atom modelling. Our method can unconditionally sample plausible backbones with over 40% validity, as measured by our custom evaluation pipeline that computes global self-consistency and local structural measurements. We also explore data augmentation protocols to address the paucity of diverse RNA structural data, which improves the novelty of designed RNA. We hope this work stimulates future research in generative models for RNA design.
@@ 2 The RNA-FrameFlow Pipeline
We are concerned with building a generative model that unconditionally outputs all-atom 3D RNA backbone structures, sampled from a distribution of realistic structures. Formally, given an RNA sequence of $N_{\text{nt}}$ nucleotides, we aim to generate a real-valued tensor $\mathbf{X}$ of shape ${N_{\text{nt}}\times 13\times 3}$ representing 3D coordinates for 13 backbone atoms per nucleotide. In the following sections, we will describe how we adapt FrameFlow (Yim et al., 2023a), an $SE(3)$ equivariant flow matching model for protein backbones, to our setting.
@@ 2.1 Representing RNA Backbones as Frames
Choice of frame atoms. We had two main considerations for selecting the atoms to create RNA frames: (1) the atoms should have roughly the same spatial orientation w.r.t. each other; and (2) the atoms should be reasonably close to the centroid in the nucleotide to reduce error accumulation when placing the furthest non-frame atoms. We choose $\{C3^{\prime},C4^{\prime},O4^{\prime}\}$ as these atoms spatially shift the least in naturally occurring RNA (Harvey & Prabhakaran, 1986). The non-frame backbone atoms – the remaining atoms in the ribose ring and the phosphate group atoms – are parameterized by torsion angles to account for their relative conformational flexibility. This choice of frame enables models to learn ring puckering, the planar rotation of the ribose ring about the $C4^{\prime}-C5^{\prime}$ bond which affects how the RNA interacts with partners to form complexes (Clay et al., 2017). We are actively evaluating alternate choices of RNA frames.
@@ 2.2 $SE(3)$ Flow Matching on RNA Frames
Input. Given a set of 3D coordinates, a simultaneous rotation and translation $(r,x)$ forms an orientation-preserving rigid-body transformation of the coordinates. The set of all such transformations in 3D is the Special Euclidean group $SE(3)$, which composes the group of 3D rotations $SO(3)$ and 3D translations in $\mathbb{R}^{3}$. We can represent an RNA frame $T=(r,x)$ as a translation $x\in\mathbb{R}^{3}$ from the global origin to place $C4^{\prime}$ and a rotation $r\in SO(3)$ to orient $C3^{\prime}-C4^{\prime}-O4^{\prime}$. Compared to working with raw 3D coordinates for each backbone atom, using the frame representation entails performing flow matching on the space of $SE(3)$. This is an inductive bias to reduce the degrees of freedom the generative model needs to learn. Instead of predicting 13 correlated 3D coordinates independently (39 quantities) for each nucleotide, we instead predict one 3D coordinate (of $C4^{\prime}$) and one 3$\times$3 rotation matrix (12 quantities). We follow Chen & Lipman (2024) and Yim et al. (2023a)’s framework for flow matching on $SE(3)$, which we summarise subsequently.
@@
[…]
@@
[…]
@@
Auxiliary losses. The inclusion of auxiliary loss terms to the objective in Equation 4 can be seen as a form of adding domain knowledge into the training process (Yim et al., 2023b). We include 3 additional losses that operate on the all-atom structure inferred from the predicted frames, weighted by tunable coefficients to modulate their contribution to the total loss:
@@
  • Coordinate MSE $\mathcal{L}_{\text{bb}}$: A direct all-atom MSE is computed between generated and ground truth coordinates. Here, $a,\hat{a}$ are the ground truth and predicted atomic coordinates for the frame atoms:
@@
  • Torsional loss $\mathcal{L}_{\text{tors}}$: An angular loss between the 8 predicted torsions by the auxiliary MLP head and the angles from the ground truth all-atom structure. Suppose $\phi\in\Phi_{n}$ and $\hat{\phi}\in\hat{\Phi}_{n}$ are the ground truth and predicted torsion angles for residue $n$, we compute:
@@
[…]
where $c=10$ is a tunable hyperparameter governing the exponential sampling schedule for rotations.
@@ 3 Experiments
3D RNA structure dataset. RNAsolo (Adamczyk et al., 2022) is a recent dataset of RNA 3D structures extracted from isolated RNAs, protein-RNA complexes, and DNA-RNA hybrids from the Protein Data Bank (as of January 5, 2024). The dataset contains 14,366 structures at resolution $\leq 4$ ($1$ = 0.1nm). We select sequences of lengths between 40 and 150 nucleotides (5,319 in total) as we envisioned this size range contains structured RNAs of interest for design tasks.
@@ 4.1 Global evaluation of generated RNA backbones
We begin by analyzing RNA-FrameFlow’s samples using the aforementioned evaluation metrics. For validity, we report percentage of samples with scTM $\geq 0.45$; for diversity, we report the ratio of unique structural clusters to total valid samples; and for novelty, we report the highest average pdbTM to a match from the PDB. For each sequence length between 40 and 150, at intervals of 10, we generate 50 backbones. Table 1 reports these metrics across different variants for the number of denoising steps $N_{T}$. The average scTM and scRMSD of valid samples are $0.641\pm 0.161$ and $2.298\pm 0.892$ respectively. We compare our model to MMDiff (Morehead et al., 2023), a protein-RNA-DNA complex co-design diffusion model. As the original best-performing version of MMDiff was trained on shorted RNA sequences, we retrain it on our training set. We also inverse-fold MMDiff’s backbones using gRNAde.
We identify $N_{T}=50$ as the best-performing model that balances validity, diversity, and novelty; furthermore, it takes 4.74 seconds (averaged over 5 runs) to sample a backbone of length 100, as opposed to 27.3 seconds for MMDiff with 100 diffusion steps. We note that increasing $N_{T}$ does not improve validity despite allowing the model to perform more updates to atomic coordinate placements. Our model also outperforms MMDiff. On manual inspection, samples from MMDiff had significant chain breaks and disconnected floating strands; see Appendix C.1.
Table 1: Unconditional RNA backbone generation. We evaluate the performance of RNA-FrameFlow for multiple values for denoising steps $N_{T}$. The best-performing model uses $N_{T}=50$ steps, taking 4.74s to sample a backbone of length 100. We also provide the average self-consistency TM-score and RMSD value for all valid samples. We green-highlight the best result per column.
Model Timesteps $N_{T}$ Validity $\uparrow$ Diversity $\uparrow$ Novelty $\downarrow$
RNA-FrameFlow $10$ 16.7 0.62 0.70
@@ 4.2 Local evaluation with structural measurements
We also show RNA Ramachandran angle plots for generated samples and the training distribution in Figure 4 (Subplot 4). Keating et al. (2011) introduced $\eta-\theta$ plots, similar to Ramachandran angle plots for proteins, that track the separate dihedral angles formed by $\{C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1},P_{i+2}\}$ and $\{P_{i},C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1}\}$ respectively, for each nucleotide $i$ along the chain. We observe that the dihedral angle distribution from RNA-FrameFlow closely recapitulates the angular distribution from naturally occurring RNA structures in the training set.
@@ 4.3 Generation quality across sequence lengths
We next investigate how sequence length affects the global realism of generated samples (measured by scTM). Figure 3 (Left) shows the performance of RNA-FrameFlow for different sequence lengths. We observe our model generates samples with high scTM for specific sequence lengths like 50, 60, 70, and 120 while generating poorer quality structures for other lengths. We believe the overrepresentation of certain lengths in the training distribution causes the fluctuation of TM-scores. We can also partially attribute this to the inherent length bias of RhoFold; see Appendix A.3. With better structure predictors, we expect more samples to be valid. We provide additional local evaluations of angular distributions in Appendix C.3.
@@ 4.4 Data preparation protocols
Due to the overrepresentation of RNA strands of certain lengths (mostly corresponding to tRNA or 5S ribosomal RNA) in our training set, our models generate close likenesses for those lengths that achieve high self-consistency but are not novel folds. To avoid this memorized recapitulation and promote increased diversity among samples, we sought to develop data preparation protocols to balance RNA folds across sequence lengths. We identically train RNA-FrameFlow on these data splits for $120$K gradient steps, with results reported in Table 3.
@@ 5 Limitations and Discussions
Physical violations. While well-trained models usually generate realistic RNA backbones, we do observe some physical violations: generated backbones sometimes have chains that are either too close by or directly clash with one another, are highly coiled, have excessive loops and unrealistically intertwined helices, or have chain breaks. We highlight these limitations in Figure 5. RNA tertiary structure folding is driven by base pairing and base stacking which influence the formation of helices, loops, and other tertiary motifs (Vicens & Kieft, 2022) . Base pairing refers to nucleotides along adjacent chains forming hydrogen bonds, while base stacking involves interactions between rings of adjacent nucleotide bases along a chain. To our knowledge, all current deep learning models operate on individual nucleotides, only implicitly learning base pairing and stacking. Developing explicit representations of these interactions as part of the architecture may further minimize physical violations and provide stronger inductive biases to learn complex tertiary RNA motifs. We analyze steric clashes in our generated backbones in Appendix C.4.
@@
We would also like to note that the models we use for structure prediction and inverse folding may be similarly biased to perform well for certain sequence lengths, leading to the overall pipeline being reliable for commonly occurring lengths and unreliable for less frequent ones (see Appendix A.3 for an analysis on RhoFold). We evaluated preliminary strategies for structural clustering and cropping augmentations during training, which improved the novelty of designed structures but led to fewer designs passing the validity filter. Overall, the relative scarcity of RNA structural data compared to proteins necessitates greater care in preparing data pipelines for scaling up training and incorporating inductive biases into generative models.
@@ 6 Related Work
Independently and concurrent to our work, Nori & Jin (2024) propose RNAFlow, an SE(3) flow matching model to co-design RNA sequence and structure conditioned on protein partners. At each denoising step, RNAFlow uses a protein-conditioned variant of gRNAde (Joshi et al., 2025) to inverse fold noised structures, followed by RoseTTAFold2NA (Baek et al., 2022a) to predict the structure of the designed sequence. The performance of RNAFlow is upper-bounded by RoseTTAFold2NA as a pre-trained structure generator, which is kept frozen and not developed for designed RNAs which do not have co-evolutionary MSA information. Our work tackles de novo 3D RNA backbone generation, an orthogonal design task of sampling RNA backbone structures. We train RNA structure generation models from scratch, akin to recent developments in protein design (Yim et al., 2023b; a; Bose et al., 2023; Lin & AlQuraishi, 2023). Backbone generation followed by inverse folding has shown experimental success in designing functional proteins (Dauparas et al., 2022; Watson et al., 2023; Ingraham et al., 2022), as the framework is flexible for including specific structural motifs and sequence constraints.
@@ A.2 Upper bound performance of the Self-consistency Pipeline
To quantify the total error accumulated in our self-consistency pipeline, and its impact on downstream validity, we study the extent to which gRNAde and RhoFold can retrieve the ground truth sequences and structures from the RNAsolo training set. To assess RhoFold’s structure prediction performance, we take all sequences of length 40 – 150 from RNAsolo, forward-fold (FF) them using RhoFold, and compute self-consistency metrics (TM-score, RMSD) by comparing them with the sequences’ associated 3D folds. To assess gRNAde’s sequence recovery performance, we inverse-fold (IF) 3D backbones from RNAsolo through gRNAde to get 16 likely sequences and pass them to RhoFold for forward-folding.
@@
RNAsolo + FF only 55.1 0.690 2.804
RNAsolo + IF + FF 43.7 0.663 3.085
RNA-FrameFlow + IF + FF (ours) 41.0 0.641 2.298

### v4

@@ 1 Introduction
Our Contributions. We develop RNA-FrameFlow, the first generative model for 3D RNA backbone design, which trains on RNA of length 40-150 from the PDB; see Figure 1. We devise RNA frames to capture the significantly larger RNA nucleotides while still supporting the modelling of all backbone atoms implicitly. Our method can unconditionally sample plausible backbones with over 40% validity, as measured by our custom evaluation pipeline that computes global self-consistency and local structural measurements. We also explore data augmentation protocols to address the paucity of diverse RNA structural data, which improves the novelty of designed RNA. We hope this work stimulates future research in generative models for RNA design.
@@ 2 The RNA-FrameFlow Pipeline
We are concerned with building a generative model that unconditionally outputs 3D RNA backbone structures, sampled from a distribution of realistic structures. Formally, given an RNA sequence of $N_{\text{nt}}$ nucleotides, we aim to generate a real-valued tensor $\mathbf{X}$ of shape $N_{\text{nt}}\times 13\times 3$} representing 3D coordinates for 13 backbone atoms per nucleotide. In the following sections, we will describe how we adapt FrameFlow (Yim et al., 2023a), an $SE(3)$ equivariant flow matching model for protein backbones, to our setting.
@@ 2.1 Representing RNA Backbones as Frames
Choice of frame atoms. We consider two main factors for selecting the atoms to create RNA frames: (1) atoms should have roughly the same spatial orientation w.r.t. each other; and (2) atoms should be reasonably close to the centroid in the nucleotide to reduce error accumulation when placing the furthest non-frame atoms. We choose $\{C3^{\prime},C4^{\prime},O4^{\prime}\}$ as these atoms have relatively lower atomic displacement in natural RNA (Harvey & Prabhakaran, 1986); see Appendix C.5 for empirical evidence on atomic displacements. The remaining non-frame backbone atoms in the ribose ring and the phosphate group are parameterized by torsion angles to account for their relative conformational flexibility. We also believe this choice of frame enables models to capture ring puckering, by which the five ribose atoms contort the ideal plane of the sugar ring to alleviate steric strain. This also influences how RNA interacts with partners to form complexes (Clay et al., 2017); see Appendix C.6 for additional details and results on puckering.
@@ 2.2 $SE(3)$ Flow Matching on RNA Frames
Input. Given a set of 3D coordinates, a simultaneous rotation and translation $(r,x)$ forms an orientation-preserving rigid-body transformation of the coordinates. The set of all such transformations in 3D is the Special Euclidean group $SE(3)$, which composes the group of 3D rotations $SO(3)$ and 3D translations in $\mathbb{R}^{3}$. We can represent an RNA frame $T=(r,x)$ as a translation $x\in\mathbb{R}^{3}$ from the global origin to place $C4^{\prime}$ and a rotation $r\in SO(3)$ to orient $C3^{\prime}-C4^{\prime}-O4^{\prime}$. Compared to working with raw 3D coordinates for each backbone atom, using the frame representation entails performing flow matching on the space of $SE(3)$. This is an inductive bias to reduce the degrees of freedom the generative model needs to learn. Instead of predicting 13 correlated 3D coordinates independently (39 quantities) for each nucleotide, we instead predict one 3D coordinate (of $C4^{\prime}$) and one $\times3$ rotation matrix (12 quantities). We follow Chen & Lipman (2024) and Yim et al. (2023a)’s framework for flow matching on $SE(3)$, which we summarise subsequently.
@@
$\text{Translations:\quad}x_{t}\ =\ tx_{1}+(1-t)x_{0}\ ,\quad\quad\quad\text{Rotations:\quad}r_{t}\ =\ \operatorname{exp}_{r_{0}}(t\cdot\operatorname{log}_{r_{0}}(r_{1}))\ .$ (2)
@@
$\text{Translations:\quad}u_{t}(x^{(n)}|x_{0}^{(n)},x_{1}^{(n)})=x_{1}^{(n)}\ ,\quad\quad\quad\text{Rotations:\quad}u_{t}(r^{(n)}|r_{0}^{(n)},r_{1}^{(n)})=\operatorname{log}_{r_{t}^{(n)}}(r_{1}^{(n)})\ .$ (3)
@@
Auxiliary losses. The inclusion of auxiliary loss terms to the objective in Equation 4 can be seen as a form of adding domain knowledge into the training process (Yim et al., 2023b). We include 3 additional losses that operate on all the backbone atoms from the predicted frames, weighted by tunable coefficients to modulate their contribution to the total loss:
@@
  • Coordinate MSE $\mathcal{L}_{\text{bb}}$: A direct MSE is computed between generated and ground truth coordinates. Here, $a,\hat{a}$ are the ground truth and predicted atomic coordinates for the frame atoms:
@@
  • Torsional loss $\mathcal{L}_{\text{tors}}$: An angular loss between the 8 predicted torsions by the auxiliary MLP head and the angles from the ground truth structure with all backbone atoms. Suppose $\phi\in\Phi_{n}$ and $\hat{\phi}\in\hat{\Phi}_{n}$ are the ground truth and predicted torsion angles for residue $n$, we compute:
@@
$\text{Translations:\quad}x_{t+\Delta t}\$$=\ x_{t}+\Delta t\cdot(\hat{x}_{1}-x_{t})\ ,$ (9)
$\text{Rotations:\quad}r_{t+\Delta t}\$$=\ \operatorname{exp}_{r_{t}}(\ c\ \Delta t\cdot\operatorname{log}_{r_{t}}(\hat{r}_{1}))\ ,$ (10)
where $c=$ is a tunable hyperparameter governing the exponential sampling schedule for rotations.
@@ 3 Experiments
3D RNA structure dataset. RNAsolo (Adamczyk et al., 2022) is a recent dataset of RNA 3D structures extracted from isolated RNAs, protein-RNA complexes, and DNA-RNA hybrids from the Protein Data Bank (as of January 5, 2024). The dataset contains 14,366 structures at resolution $\leq 4$ Å ($1$ Å = 0.1nm). We select sequences of lengths between 40 and 150 nucleotides (5,319 in total) as we envisioned this size range contains structured RNAs of interest for design tasks.
@@ 4.1 Global evaluation of generated RNA backbones
We begin by analyzing RNA-FrameFlow’s samples using the aforementioned evaluation metrics. For validity, we report percentage of samples with scTM $\geq 0.45$; for diversity, we report the ratio of unique structural clusters to total valid samples; and for novelty, we report the highest average pdbTM to a match from the PDB. For each sequence length between 40 and 150, at intervals of 10, we generate 50 backbones. Table 1 reports these metrics across different variants for the number of denoising steps $N_{T}$. The average scTM and scRMSD of valid samples are $0.641\pm 0.161$ and $2.298\pm 0.892$ respectively. We compare our model to MMDiff (Morehead et al., 2023), a protein-RNA-DNA complex co-design diffusion model. As the original best-performing version of MMDiff was trained on shorter RNA sequences, we retrain it on our training set. We also inverse-fold MMDiff’s backbones using gRNAde.
We identify $N_{T}=50$ as the best-performing model that balances validity, diversity, and novelty; furthermore, it takes 4.74 seconds (averaged over 5 runs) to sample a backbone of length 100, as opposed to 27.3 seconds for MMDiff with 100 diffusion steps. We note that increasing $N_{T}$ does not improve validity despite allowing the model to perform more updates to atomic coordinate placements. Our model also outperforms MMDiff. On manual inspection, samples from MMDiff had significant chain breaks and disconnected floating strands; see Appendix C.1. We report fine-grained self-consistency metrics factoring in the rotational component of frames as well as all backbone atoms in Appendix B.4.
Table 1: Unconditional RNA backbone generation. We evaluate the performance of RNA-FrameFlow for multiple values for denoising steps $N_{T}$. The best-performing model uses $N_{T}=50$ steps, taking 4.74s to sample a backbone of length 100. We include the forward-folding model used in the self-consistency pipeline. We green-highlight the best result per column.
Model Timesteps $N_{T}$ % Validity $\uparrow$ Diversity $\uparrow$ Novelty $\downarrow$
[…]
@@ 4.2 Local evaluation with structural measurements
We also show RNA Ramachandran angle plots for generated samples and the training distribution in Figure 4 (Subplot 4). Keating et al. (2011) introduce $\eta-\theta$ plots, similar to Ramachandran angle plots for proteins, that track the separate dihedral angles formed by $\{C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1},P_{i+2}\}$ and $\{P_{i},C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1}\}$ respectively, for each nucleotide $i$ along the chain. We observe that the dihedral angle distribution from RNA-FrameFlow closely recapitulates the angular distribution from naturally occurring RNA structures in the training set. We provide additional evidence on the successful modeling of ring puckering in RNA-FrameFlow’s generated backbones in Appendix C.6.
@@ 4.3 Generation quality across sequence lengths
We next investigate how sequence length affects the global realism of generated samples (measured by scTM). Figure 3 (Left) shows the performance of RNA-FrameFlow for different sequence lengths. We observe our model generates samples with high scTM for specific sequence lengths like 50, 60, 70, and 120 while generating poorer quality structures for other lengths. We believe the overrepresentation of certain lengths in the training distribution causes the fluctuation of TM-scores. We can also partially attribute this to the inherent length bias of RhoFold; see Appendix A.3. With better structure predictors, we expect more samples to be valid. We provide additional local evaluations of angular distributions in Appendix C.3. We also provide an ablation using Chai-1 (Boitreaud et al., 2024) as the forward-folding model in Appendix B.3; we do not observe a significant change in validity with Chai-1.
@@ 4.4 Data preparation protocols
Due to the overrepresentation of RNA strands of certain lengths (mostly corresponding to tRNA or 5S ribosomal RNA) in our training set, our models generate close likenesses for those lengths that achieve high self-consistency but are not novel folds. To avoid this memorized recapitulation and promote increased diversity among samples, we sought to develop data preparation protocols to balance RNA folds across sequence lengths. We identically train RNA-FrameFlow on these data splits for 120120K gradient steps, with results reported in Table 3.
@@ 5 Limitations and Discussions
Physical violations. While well-trained models usually generate realistic RNA backbones, we do observe some physical violations: generated backbones sometimes have chains that are either too close by or directly clash with one another, are highly coiled, have excessive loops and unrealistically intertwined helices, or have chain breaks. We highlight these limitations in Figure 5 and analyze steric clashes in our generated backbones in Appendix C.4. RNA tertiary structure folding is driven by base pairing and base stacking which influence the formation of helices, loops, and other tertiary motifs (Vicens & Kieft, 2022). Base pairing refers to nucleotides along adjacent chains forming hydrogen bonds, while base stacking involves interactions between rings of adjacent nucleotide bases along a chain. RNA-FrameFlow attempts to model these phenomena using auxiliary losses but still operates at the backbone-only granularity, not including base atoms (barring $N1/N9$). An array of complex tertiary folds (e.g., kissing loops, tetraloops, junctions) occur through base-mediated interactions, meaning RNA-FrameFlow can only implicitly learn base pairing and stacking. Developing explicit representations of these interactions as part of the architecture may further minimize physical violations and provide stronger inductive biases to learn complex tertiary RNA motifs.
@@
We would also like to note that the models we use for structure prediction and inverse folding may be similarly biased to perform well for certain sequence lengths, leading to the overall pipeline being reliable for commonly occurring lengths and unreliable for less frequent ones. We also provide an empirical upper bound on the error accumulated in the self-consistency pipeline in Appendix A.2. We evaluated preliminary strategies for structural clustering and cropping augmentations during training, which improved the novelty of designed structures but led to fewer designs passing the validity filter. Overall, the relative scarcity of RNA structural data compared to proteins necessitates greater care in preparing data pipelines for scaling up training and incorporating inductive biases into generative models.
@@ 6 Related Work
Independently and concurrent to our work, Nori & Jin (2024) propose RNAFlow, an SE(3) flow matching model to co-design RNA sequence and structure conditioned on protein partners. At each denoising step, RNAFlow uses a protein-conditioned variant of gRNAde (Joshi et al., 2025) to inverse fold noised structures, followed by RoseTTAFold2NA (Baek et al., 2022a) to predict the structure of the designed sequence. The performance of RNAFlow is upper-bounded by RoseTTAFold2NA as a pre-trained structure generator, which is kept frozen and not developed for designed RNAs which do not have co-evolutionary MSA information. Our work tackles de novo 3D RNA backbone generation, an orthogonal design task of sampling RNA backbone structures. We train RNA structure generation models from scratch, akin to recent developments in protein design (Yim et al., 2023b; a; Bose et al., 2023; Lin & AlQuraishi, 2023). Backbone generation followed by inverse folding has shown experimental success in designing functional proteins (Dauparas et al., 2022; Watson et al., 2023; Ingraham et al., 2022), as the framework is flexible for including specific structural motifs and sequence constraints. Since the original release of RNA-FrameFlow and its source code in 2024, several methods employ a similar $SE(3)$ flow matching setup. Rubin et al. (2025) incorporate an additional sequence track for RNA structure-sequence co-design, using the MultiFlow framework from Campbell et al. (2024). Tarafder & Bhattacharya (2025) condition the 3D structure generation on pre-determined sequences and secondary structural contact maps, enabling all-atom resolution with base identities.
@@ A.2 Upper bound performance of the Self-consistency Pipeline
To quantify the total error accumulated in our self-consistency pipeline, and its impact on downstream validity, we study the extent to which gRNAde and RhoFold can retrieve the ground truth sequences and structures from the RNAsolo training set. To assess RhoFold’s structure prediction performance, we take all ground truth sequences of length 40 – 150 from RNAsolo, forward-fold (FF) them using RhoFold, and compute self-consistency metrics (TM-score, RMSD) by comparing them with the sequences’ associated 3D folds. To assess gRNAde’s sequence recovery performance, we inverse-fold (IF) 3D backbones from RNAsolo through gRNAde to get 16 likely sequences and pass them to RhoFold for forward-folding.
@@
RNAsolo + $F$ only
55.1 0.690 2.804
RNAsolo + IF + $F$
43.7 0.663 3.085
RNA-FrameFlow + IF + $F$ (ours)
41.0 0.641 2.298